OpenStudy (anonymous):

7-x^3(1/2x) when x=-2

OpenStudy (anonymous):

Replace x by -2 and the solve..

OpenStudy (anonymous):

\[7 - (-2)^3(\frac{1}{2}(-2))\]

OpenStudy (anonymous):

I keep getting 5 but my calculator says its -1

OpenStudy (anonymous):

Is x in the denominator @J.L.

OpenStudy (anonymous):

Is x in the denominator ??

OpenStudy (anonymous):

the answer is 5

OpenStudy (anonymous):

@waterineyes , the x is the denominator, obviously

OpenStudy (anonymous):

I think @J.L. can also say this..

OpenStudy (anonymous):

@J.L., the reason your calculator gives you -1, because it considers it's 1/2 * -2, and not 1/(2*-2)

OpenStudy (anonymous):

\[7 - (-2)(4) \times \frac{1}{(2) (-2)} \implies 7 - 2 = 5\]